The difference between two positive integers is 36, one integer is three times as great as the other.
Answer by stanbon(57357)

The difference between two positive integers is 36, one integer is three times as great as the other.
-------------------------------------
x = 3y
x-y = 36
----
Substitute for "x" and solve for "y":
3y-y = 36
2y = 36
y = 18
-------
x = 3y = 3*18 = 54
---
Solution:
(54,18)
==================
Cheers,
stan H.